# Clustering and Data Augmentation to Improve Accuracy of Sleep Assessment and Sleep Individuality Analysis

### v1

@@ 1 Introduction
In the previous studies [8, 9], the feature representation of sleep sound events was obtained using Variational AutoEncoder (VAE) [10], and the classification of sleep quality was based on Long Short-Term Memory (LSTM) [11]. In this study, we have extended these methods. The novelty and advantages of the proposed approach are as follows:
  • Clustering of sleep sound events using Gaussian Mixture Model (GMM) makes it easier to interpret sleep sound events by humans.
  • By using the membership probabilities of clusters as representations of sleep sound events, the dimensionality of LSTM is reduced, allowing for an increase in the amount of training data through the proposed data augmentation. As a result, the generalization performance of the model improves.
  • With TimeSHAP [12], it becomes possible to analyze important sleep sound events for sleep quality classification and individual-specific features.
@@ 4.1 Dataset
In this study, data collected with the cooperation of participants from a wide range of age groups was utilized. The data collection was conducted with the approval of the Ethics Committee of SANKEN, Osaka University. Participants were asked to record their sleep sounds in their private bedrooms at home for one month. Recording was done using a smartphone (Zenfone Live Android 7.0). Additionally, participants filled out questionnaires. Before bedtime, they provided answers regarding their physical and mental fatigue, presence of illnesses or injuries, etc. After waking up, they rated their sleep satisfaction and the indoor environment during sleep. Sleep satisfaction was rated on a five-point scale: “very satisfied,” “satisfied,” “neutral,” “unsatisfied,” and “very unsatisfied.” Moreover, participants’ profiles such as age and gender were recorded.
@@
Table 1: Subjects’ basic information
subject ID age gender # satisfied days # unsatisfied days
1 43 F 8 10
2 23 M 12 8
3 21 F 10 11
@@ 4.2 Classification Result of Sleep Satisfaction
In this experiment, we conducted sleep quality classification using the proposed method and examined its classification accuracy. First, we trained the LSTM with varying conditions, the dimensionality of VAE latent space, and the number of clusters in GMM. The LSTM had a single hidden layer with 50 nodes, and a dropout rate of 0.2 was applied between the hidden and output layers. The output layer utilized a Sigmoid activation function. We employed the Adam optimizer [23] with a learning rate of 0.001. The sampling size of the sound events used in LSTM training was set according to the number of events for each participant. The multiplying factor of the augmentation was set to 5,000 for all participants. To prevent overfitting, we employed Early Stopping if the loss on the test data did not improve for more than 5 epochs. To increase the number of trials, we performed 4-Fold Cross Validation five times with different random seeds to determine the optimal conditions for each participant. The results are presented in Tables 2, 3, and 4. It was observed that the optimal conditions for the dimensionality of VAE latent space and the number of clusters for achieving high accuracy varied among participants.
@@
None 0.800±0.044 0.600±0.00 0.717±0.066
200 0.845±0.075 0.740±0.102 0.858±0.068
500 0.818±0.053 0.730±0.129 0.827±0.076
1,000 0.845±0.075 0.780±0.081 0.867±0.100
2,000 0.903±0.040 0.840±0.073 0.893±0.048
5,000 0.893±0.036 0.826±0.061 0.948±0.050
@@ 4.4 Comparison with the Conventional Method
To ensure approximately equal computational time per epoch between the proposed and conventional methods, we adjusted the data augmentation factor for the conventional method. We then compared the accuracy of sleep quality assessment for each method. Additionally, we conducted a one-sided t-test with a significance level of 5% to confirm whether there was a significant difference in the mean values between the proposed and conventional methods.
The experimental results are presented in Table 6. For the conventional method, the latent space dimension of VAE and the multiplying factor of data augmentation were (100, 120) for Subject 1, (20, 300) for Subject 2, and (150, 200) for Subject 3, respectively. The values in the table for the conventional method were obtained using the optimal latent space dimension of VAE. It can be observed that the proposed method outperforms the conventional method in terms of accuracy for all subjects. Furthermore, the accuracy of the proposed method was significantly higher than that of the conventional method for Subject 1 and Subject 2. Therefore, it can be concluded that the proposed method achieves at least the same level of accuracy as the conventional method when compared.
@@ 5 Conclusion
This study proposed an interpretable method for sleep quality classification using clustering of latent representations of sleep sound events. The experimental results demonstrated that the proposed method achieves high accuracy in classifying sleep satisfaction for a night. Furthermore, we showed significant differences in the importance of clusters based on sleep satisfaction, the ability of the proposed method to analyze individual sleep characteristics, and identify areas for improvement. Additionally, we demonstrated that the proposed method achieves equal or higher accuracy to the conventional method.
The following are the future challenges. The first challenge lies in the significant variation of optimal VAE dimensions and cluster numbers among individuals. Efficient tuning methods will be necessary when deploying sleep evaluation applications using the proposed method. The second challenge involves the manual labeling of sleep sound events. Since there are individual differences of the sounds even for the same events, it is necessary to label sleep sound events for each individual. Manual labeling becomes difficult when deploying on a large scale. Therefore, there is a need to streamline labeling through methods such as constructing an event classifier and utilizing transfer learning.

### v2

@@ 1 Introduction
In the previous studies [8, 9], the feature representation of sleep sound events was obtained using Variational AutoEncoder (VAE) [10], and the classification of sleep quality was based on Long Short-Term Memory (LSTM) [11]. In this study, we extended a classification method based on VAE-LSTM by adding explainability to the classification of sleep quality. To provide explainability, it is necessary to identify the types of sound events; however, labeling each event individually is challenging. Therefore, we introduced sound event clustering to facilitate the labeling of sound events. Additionally, we proposed a method to obtain explanations for the trained classification model using TimeSHAP[12]. This method can present the contribution of each sound event cluster to the classification. Through an experiment with three subjects conducted in a home environment over approximately 20 days, we demonstrated that it is possible to analyze which types of events are involved in determining sleep quality and to identify the unique sleep characteristics of each subject.
Moreover, while conventional methods used discrete power spectra of sound events as feature vectors for input to the VAE, in this study, we proposed a method that uses the probability of belonging to each cluster as input. By combining this with data augmentation, an improvement trend in classification accuracy was observed in experiments with the three subjects.
The novelty of the proposed method compared to the conventional method of sleep quality classification using VAE-LSTM is as follows:
  • We proposed a framework that introduces clustering of sleep sound events, making it possible to evaluate the types of sound events that affect sleep quality classification by LSTM.
  • We proposed a method that represents sleep sound events using the probability of belonging to each cluster. Through experiments, we demonstrated that combining this with data augmentation could lead to improved accuracy in sleep quality classification.
@@ 4.1 Dataset
In this study, data collected with the cooperation of participants from a wide range of age groups was utilized. The data collection was conducted with the approval of the Ethics Committee of SANKEN, Osaka University. Participants were asked to record their sleep sounds in their private bedrooms at home for one month. Recording was done using a smartphone (Zenfone Live Android 7.0). The sampling frequency of the microphone is 48 kHz. Additionally, participants filled out questionnaires. Before bedtime, they provided answers regarding their physical and mental fatigue, presence of illnesses or injuries, etc. After waking up, they rated their sleep satisfaction and the indoor environment during sleep. Sleep satisfaction was rated on a five-point scale: “very satisfied,” “satisfied,” “neutral,” “unsatisfied,” and “very unsatisfied.” Moreover, participants’ profiles such as age and gender were recorded.
@@
Table 1: Subjects’ basic information
subject ID age gender # satisfied days # unsatisfied days total rec. time
1 43 F 8 10 96h
2 23 M 12 8 125.5h
3 21 F 10 11 147.5h
@@ 4.2 Classification Result of Sleep Satisfaction
In this experiment, we conducted sleep quality classification using the proposed method and examined its classification accuracy. First, we trained the LSTM with varying conditions, the dimensionality of VAE latent space, and the number of clusters in GMM. The LSTM had a single hidden layer with 50 nodes, and a dropout rate of 0.2 was applied between the hidden and output layers. The output layer utilized a Sigmoid activation function. We employed the Adam optimizer [23] with a learning rate of 0.001. And a binary cross-entropy is used for the loss function. The sampling size of the sound events used in LSTM training was set according to the number of events for each participant. The multiplying factor of the augmentation was set to 5,000 for all participants. To increase the number of trials, we performed 4-Fold Cross Validation five times with different random seeds to determine the optimal conditions for each participant. The training was terminated after 10 epochs, as the loss on the training data had generally converged. Table 2 shows the number of sleep sound events used for the training data (before data augmentation) and test data for each subject.
Table 2: The number of sleep sound events for each subject. It varies due to the different number of days in the cross-validation splits.
ID Training data (w/o data augmentation) Test data
1 5,200 (13 days) / 5,600 (14 days) 2,000 (5 days) / 1,600 (4 days)
2 3,000 (15 days) 1,000 (5 days)
3 7,500 (15 days) / 8,000 (16 days) 3,000 (6 days) / 2,500 (5 days)
The results are presented in Tables 3, 4, and 5. Hereafter, all “$\pm$” in the tables indicate the standard deviation. It was observed that the optimal conditions for the dimensionality of VAE latent space and the number of clusters for achieving high accuracy varied among participants.
@@
None 0.800 $\pm$ 0.044 0.600$\pm$0.00 0.717$\pm$0.066
200 0.845$\pm$0.075 0.740$\pm$0.102 0.858$\pm$0.068
500 0.818$\pm$0.053 0.730$\pm$0.129 0.827$\pm$0.076
1,000 0.845$\pm$0.075 0.780$\pm$0.081 0.867$\pm$0.100
2,000 0.903$\pm$0.040 0.840$\pm$0.073 0.893$\pm$0.048
5,000 0.893$\pm$0.036 0.826$\pm$0.061 0.948$\pm$0.050
Here, for subjects 1 and 2, performance peaked at 2,000x data augmentation. In this data augmentation process, increasing the augmentation rate did not always lead to improved performance. While the accuracy on the training data was between 0.88 and 0.91 at 2,000x augmentation, it increased to between 0.96 and 0.98 at 5,000x. However, since this data augmentation method generates new event sequences by sampling sound event sequences, excessive data augmentation may lead to overfitting.
@@ 4.4 Comparison with the Conventional Method
To ensure approximately equal computational time per epoch between the proposed and conventional methods, we adjusted the data augmentation factor for the conventional method. We then compared the accuracy of sleep quality assessment for each method. The data augmentation rate was set to 2,000x for subjects 1 and 2 and 5,000x for subject 3, for both the proposed method and the conventional method. Additionally, we conducted a one-sided t-test with a significance level of 5% to confirm whether there was a significant difference in the mean values between the proposed and conventional methods.
The experimental results are presented in Table 7. For the conventional method, the latent space dimension of VAE and the multiplying factor of data augmentation were (100, 120) for Subject 1, (20, 300) for Subject 2, and (150, 200) for Subject 3, respectively. The values in the table for the conventional method were obtained using the optimal latent space dimension of VAE. For subjects 1 and 3, the proposed method achieved a statistically significantly higher mean accuracy compared to the conventional method, while for subject 2, the results were comparable. Therefore, it can be concluded that the proposed method achieves at least the same level of accuracy as the conventional method when compared.
@@ 5 Conclusion
In this study, we proposed a novel method for sleep quality classification over a single night using VAE and LSTM by introducing clustering of sleep sound events and using TimeSHAP to identify the types of sound events involved in the classification as explanations. Through an experiment with three subjects conducted in a home environment over approximately 20 days, we demonstrated that there is a significant difference in the importance of clusters depending on sleep satisfaction, and we analyzed each subject’s unique sleep characteristics, suggesting possible areas for improvement. Additionally, we proposed a method that uses the probability of belonging to each cluster as a representation of sound events. By combining this with data augmentation of event sequences, we observed a trend toward improved classification accuracy compared to the conventional method in the experiments with the three subjects.
The future challenges are as follows: The first challenge is that the number of subjects in this study was limited to three, so it will be necessary to verify the generalizability of the method with a larger number of subjects.
The second challenge is that the optimal dimensionality of the VAE and the number of clusters vary significantly between individuals. When deploying applications for sleep evaluation using the proposed method, an efficient tuning approach will be needed.
The third challenge is the manual labeling of sleep sound events. Since there are individual differences in sleep sound events, even for the same phenomena, it is necessary to label sound events for each person. Manual labeling would be difficult when scaling up to a large number of users. Therefore, it will be necessary to streamline the labeling process by training event classifiers or using transfer learning.